OpenStudy (redbubbles):

what is the volume of the pyramid to the nearest whole unit? the height is 7yd the width is 9yd and the length is 7yd the answer choices are A147yd^3 B 175 yd^3 C 221 yd^3 D 441 yd^3 I think its B can someone help?

OpenStudy (anonymous):

\[V(Pyramid) = \frac{ lwh }{ 3 } = \frac{ 7(9)(7) }{ 3 }\]

OpenStudy (anonymous):

\[= 7(3)(7) = ?\]

OpenStudy (redbubbles):

so it would be 147 yd^3